my question is this?
1.In a box of 12 parts 5 are defective. Two parts are selected with out replacement. What is the probability that a least one of the parts is defective?
2. A pizza can be made with 7 different cheese and 2 of 8 different topping. how many different combinations are possible.
i want to know how to solve this problem? step-by-step leading to the answer.

Answer by edjones(8007)

1)
nCr=combination of n things taken r at a time=n!/((n-r)!*r!)
Probability none are defective: 7C2/12C2 = 21/66 = 7/22
Probability none at least one is defective: 1- 7/22 = 15/22
.
Ed